Alert, 2 May, 2006
Received from Trine Dahl-Jensen
Edited by Henrik Højmark Thomsen, GEUS
Finally – the last seismic data is retrieved
1 May, 2006
Weather in Alert: -18, sunny, calm.
Finally we got a change in the terrible weather, which has made so much trouble for us the last 10 days. The weather was good this morning with only a few fog patches over the ice. It looked like the day, we should make an attempt to do the reduced programme on the last seismic line – the so called cross line running east-west perpendicular to the two other lines, we have measured along. Our original plan was to detonate 11 shots along the 200 km long line and record the shots with 150 instruments, which should be deployed on the ice along the line. However the many days with no flying have forced us only to go for a reduced programme on the line. We now plan to detonate six shots along the mid 100 km of the line and record the signals with 72 receivers.
During all the weather days in Alert we had prepared our self for this moment. Explosives were already in the cache on the cross line and at 9:00 in the morning all three helicopters took off from Alert heading for the line. During the next hours all six shots were loaded by the three helicopter teams. Meanwhile the Twin Otter was flying out all the 72 receivers to the cache, and later the helicopter teams started deploying the instruments along the line. After the deployment two teams visited the shot sites and blasted all six shots and short after all three teams were busy recovering the instruments. The operation went well and we succeeded to pickup all the 72 receivers, but the pickup of the westernmost instruments was tricky due to fog patches over this part of the line.
We returned tired, but very happy back to Alert at 23:30 after a very long working day. Finally we had retrieved the data from the last seismic line, and the seismic programme is now finished. It was a very good day.
2 May, 2006
Weather in Alert: -17, sunny, light wind.
Back to normal. Another weather day in Alert with no flying. The weather was fine in Alert but there were fog and low cloud over our working area on the ice. We were obviously very lucky with the few hours of flying weather yesterday. The BG-team has waited the whole day long for improvements. They are eager to continue their measurements of water depths and gravity.
All instruments were emptied for data and we have started packing the seismic equipment in preparation for a transport flight from Alert to Edmonton Saturday May 6.
We are crossing our fingers for better weather, so that the BG-team can continue their work. Our plan is, that they will work in two teams – each with a helicopter. And if the weather is really fine, we will make a final attempt to recover the five instruments on the outer line together with the satellite beacon at position OS7, which was left there to help the search for the receivers. They have now been standing on the ice since the operation on 21 April, where we could not find them.